# Large-scale peculiar velocities in the universe

### v2

@@ 2.4.2 Resolution of historical conflicts
Another significant resolution concerns the apparent conflict between peculiar velocity surveys and other velocity-sensitive probes. An analysis by Boruah et al [147] showed that, once sparse sampling, survey selection, and cosmic variance are properly accounted for within consistent modeling frameworks, independent probes of large-scale motions yield mutually consistent constraints on the velocity field. This reconciliation has been particularly important for establishing the reliability of bulk-flow measurements across different methodologies.
@@ 3.1 Fundamental Principles of Peculiar Velocity Measurement
Two primary methodological approaches have emerged for characterizing bulk flows. The first involves measuring individual galaxy peculiar velocities and then calculating their weighted average within specified volumes. The second employs likelihood methods to fit parametric models directly to the observed velocity field [7, 11].
@@ 3.3.2 Novel techniques and emerging methods
where is the Hubble constant, and is the speed of light.
@@ 4 Theoretical framework
This section outlines the basics of the 1+3 covariant approach to relativistic cosmology. The formalism. which dates back to the 1950s (e.g. see [199, 200]), has since been used in numerous studies by many authors (see [88, 89] for recent reviews with more details and references).
@@ 4.1 Covariant spacetime splitting
The covariant approach utilises a local 1+3 splitting of the spacetime into time and 3-dimensional space. This is achieved by introducing a group of observers, “living” along timelike worldlines with their associated 3-spaces orthogonal to them. Instead of using the metric tensor, the covariant formalism employs the Ricci and the Bianchi identities, together with the Einstein field equations [88, 89].
Consider a group of observers with 4-velocity , where are their (generalised) coordinates and their proper time, so that .111We adopt spacetimes with metric () of signature () and use geometrised units with , unless stated otherwise. Also, Latin indices take their values between and , while their Greek counterparts run from to . The -field determines the observers’ temporal direction and introduces a local 1+3 “threading” of the spacetime into time and 3-D space. Then, the tensor projects into the observers’ instantaneous rest space and also acts as the 3-D metric. Note that the projector is symmetric, spacelike and also satisfies the constraints and .
@@ 4.1.1 The gravitational field
respectively. Note that has a Newtonian analogue but has not (e.g. see [201]). For this reason the magnetic Weyl tensor is primarily associated with gravitational waves, although both Weyl fields are necessary for the propagation of gravitational waves. Also, both and are spacelike, symmetric and trace-free tensors.
@@ 4.2.1 Kinematics
Here, is the volume scalar, is the shear, is the vorticity and is the 4-acceleration. The last three are purely spacelike, since . The volume scalar describes expansion/contraction when positive/negative respectively. In addition, is used to define the cosmological scale factor () by means of . Given its trace-free nature, the shear monitors shape distortions under constant volume, while antisymmetric vorticity tensor describes rotation and also defines the rotational axis via the vorticity vector . Finally, nonzero 4-acceleration indicates the presence of non-gravitational forces and ensures that the observers’ worldlines are not geodesics.
@@ 4.3 The Friedmann universes
Our discussion so far refers to cosmological models that are inhomogeneous and anisotropic. Nevertheless, observations and our theoretical prejudice (the CMB and the Copernican principle) support the homogeneous and isotropic Friedmann-Robertson-Walker (FRW) models.
4.3.1 The FRW metric
@@
4.3.2 FRW cosmologies
The high symmetry of the Friedmann models ensures that time-dependent scalars are the only surviving variables. Hence, in covariant terms, the FRW cosmologies have , where is the familiar Hubble parameter, and . In addition, the isotropy of the Friedmann universes ensures that they can only accommodate perfect fluids, with and ). Finally, due to spatial homogeneity, all the 3-gradients (e.g. , , etc) vanish by default. All this means that the only nontrivial equations are the FRW versions of Raychaudhuri’s formula, together with the continuity and the Friedmann equations. These follow from (4.2.3), (4.2.11a) and (4.2.16) and are given by
@@ 4.3.2 FRW cosmologies
In FRW universes with conventional matter, effectively coincides with the particle horizon (i.e. ), in which case the Hubble radius determines the regions of causal contact as well.
@@
Consequently, always for , with as and when . On the other hand, for and when in closed models.
@@ 4.3.3 Luminosity distance
The luminosity distance of a source at redshift is , with and being the present values of the scale factor and of the source’s radial distance (e.g. see [208]). The latter follows by integrating the line element of a null geodesic. In a spatially flat FRW model, the integration gives
@@
Alternative expressions of the luminosity distance, as well as generalisations of that incorporate the (linear) effects of peculiar motions, can be found in § 10.1.1.
4.3.4 Scale-factor evolution in FRW cosmologies
In the following sections, the host universe will be typically assumed to be a spatially flat FRW model with . Nevertheless, there will be cases where the background spacetime will be allowed to have nonzero spatial curvature (see § 6.3.5 and § 7.1.8 below). It will therefore help to briefly go through the basic evolution laws of Friedmann universes with as well.
@@ 4.3.4 Scale-factor evolution in FRW cosmologies
The case: In FRW spacetimes with curved spatial sections it helps to parametrise the solutions in terms of the conformal time ( – defined so that ). Then, for , and relations (4.3.3) and (4.3.4) combine to the parametric solution
@@ 5 Cosmologies with a peculiar-velocity “tilt”
Peculiar velocities are defined and measured with respect to the reference frame of the universe, which has been typically identified with the rest-frame of the cosmic microwave photons. By definition, the latter is the only coordinate system where the CMB dipole vanishes. Nevertheless, only idealised observers follow the cosmic microwave frame. All real observers, namely those living in typical galaxies like our Milky Way, have some finite peculiar velocity with respect to the CMB. It is therefore clear that the study of peculiar motions requires tilted cosmologies, equipped with two (at least) reference frames moving relative to each other.
@@ 5.1 Relatively moving observers
Consider a tilted spacetime and allow for a family of 4-velocity fields in relative motion. Suppose also that is the reference 4-velocity of the CMB frame and is that of the -th fluid component, with . Then, the symmetric 3-tensors projecting orthogonal to and are
respectively. By construction, the above projectors satisfy the constraints , , and .
@@
One can recast the velocity boost (5.1.4) in terms of the hyperbolic “tilt” angle () between the two 4-velocity vectors (see Fig. 1). Indeed, given that and that , where , expression (5.1.4) recasts into [211]
@@ 5.3 Linear relations between the frames
Consider an FRW background universe, where the only non-vanishing variables are scalars that depend solely on time. These are the density (), the isotropic pressure (), the volume expansion (, with ) and the spatial Ricci scalar (, with and – see § 4.3.2 earlier). Then, when dealing with non-relativistic peculiar velocities (with and ), the nonlinear relations (5.2.4)-(5.2.7) linearise to
@@
respectively. Here, is the local volume scalar, with since the peculiar flow can expand or contract locally, is the local votricity vector and is the local shear tensor of the peculiar motion. The above follow from the spatial gradient of the peculiar-velocity field, which decomposes as [212]
@@ 6 Linear peculiar velocities
Bulk peculiar flows are commonplace in the universe, confirmed by several surveys over many years (e.g. see [213] for the Planck 2013/2015 constraints). Such large-scale motions are believed to have started as weak peculiar-velocity perturbations around the time of recombination, which then grew to the observed sizes and speeds driven by structure formation. There are open issues, however. The reported velocities generally lie within CDM expectations on relatively small scales, roughly Mpc [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23], but appear in excess of the typical CDM predictions on larger scales [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34]. In this section we will review the theoretical studies of peculiar velocities, referring the reader to § 9 for the observational surveys and reports.
@@ 6.1.1 Newtonian peculiar kinematics
where is the peculiar velocity of the -field relative to the (reference) -frame.555Typical Newtonian studies of peculiar motions use physical () and comoving () coordinates, with . The time derivative of the latter leads to , where , and are the total, the Hubble and the peculiar velocities respectively. On an FRW background, the above velocity relation is equivalent to the one seen in Eq. (6.1.1). The irreducible kinematics of the -field are given by the decomposition
@@
with , and respectively representing the volume scalar, the shear tensor and the vorticity of the peculiar flow.666Comparing (6.1.2) to its relativistic counterpart (see decomposition (4.2.1) in § 4.2.1), one notices the absence of an acceleration term from the right-hand sides of the Newtonian expression. This reflects the fact that, in Newtonian physics, space and time are separate and independent entities. As before (see Eq. (5.3.3) in § 5.3), positive/negative values for indicate (locally) expanding/contracting peculiar flows. Also, and are respectively the shear and vorticity of the peculiar-velocity field. Starting from (6.1.1), one arrives at the following (nonlinear) relations
@@ 6.1.2 Linear sources of peculiar velocities
In a perturbed almost-FRW Newtonian universe (filled with a pressureless perfect fluid – baryonic or/and CDM – with ), the peculiar-velocity field is a linear perturbation. Then, by identifying the -field with the CMB frame, we may set at the linear level, in which case expression (6.1.5) recasts to
According to the above, the sole source of linear peculiar velocities is the acceleration, which in the absence of pressure is given by . As a result, Eq. (6.1.6) reads
@@ 6.2.1 The quasi-Newtonian setup
The quasi-Newtonian approach is used to obtain Newtonian-like equations and solutions, by adopting a reference frame with zero linear vorticity and shear (i.e. by default). This, however, imposes strict restrictions on the spacetime and compromises its relativistic nature. For instance, in the quasi-Newtonian models, the gravitational waves are switched off as well. Perturbed relativistic spacetimes without these features are rather unnatural. One should therefore be wary before adopting the quasi-Newtonian approximation in relativistic cosmological studies (see § 6.8.2 in [89] for related discussion and “warning” comments).
The downside of employing the quasi-Newtonian approach to study cosmological peculiar velocities stems form the fact that the (purely relativistic) gravitational input of the peculiar flux (see § 5.3 earlier) is not accounted for. The direct consequence of this is that the driving force of the linear peculiar-velocity field, namely the 4-acceleration, is given by the gradient of an ad hoc (Newtonian-like) scalar potential, the evolution of which follows from an ansatz. Although this may simplify the calculations, it compromises the relativistic nature of the treatment and blurs the physics. As a result, the quasi-Newtonian analysis (inadvertently) reduces to Newtonian (see § 6.1 previously and § 6.3 below). Nevertheless, as will show in § 6.3, there is no need for any of the aforementioned severe mathematical constraints when studying linear peculiar motions relativistically, which makes the related quasi-Newtonian studies an interesting mathematical exercise but without any physical input.
@@ 6.2.2 Linear quasi-Newtonian approach
The starting point is a tilted, perturbed Einstein-de Sitter background universe, with two families of observers in relative motion. These are the idealised (reference) CMB observers with 4-velocity and their realistic (tilted) counterparts with 4-velocity . The latter are assumed to move with (non-relativistic) peculiar velocity with respect to the former, so that (see Fig. 2).777In the background spacetime, the peculiar-velocity field vanishes by default and the two 4-velocities coincide. This makes a gauge-invariant linear perturbation [215]. Assuming that the pressureless matter follows timelike geodesics, one may set in the tilted frame. Then, in line with (5.3.1b) and (5.3.4c), we have
@@
There is a serious downside, however, because is an ad hoc potential with no expression to describe it. In addition, the time evolution of the quasi-Newtonian potential is not uniquely determined and follows from the ansatz , which requires setting the shear to zero as well [41]. Moreover, when it comes to the study of peculiar velocities, the key disadvantage is that the 4-acceleration adopted in Eq. (6.2.2) does not account for the gravitational input of the peculiar flux and this severely compromises the relativistic nature of the analysis. After all this, the linear evolution of the quasi-Newtonian 4-acceleration is given by [41, 42]
@@
Before closing this section, we should point out that the general problems of the quasi-Newtonian studies, namely of adopting a reference frame with zero linear shear and vorticity (and the rest of the constraints that follow), have been known and they are clearly stated in § 6.8.2 of [89]. However, the extent and the depth of the problem was not realised because there was no direct comparison with a proper relativistic treatment that did not employ the quasi-Newtonian frame. Next, we will provide such a comparison by avoiding the quasi-Newtonian zero vorticity and shear restrictions, when studying linear peculiar velocities. In the process, we will also identify the pivotal role of the peculiar flux and show how deeply damaging it can be to bypass it. Without accounting for the gravitational input of the flux, peculiar-motion studies that may appear relativistic reduce to Newtonian.888In the literature there additional examples of studies involving peculiar motions, which start relativistically but reduce to Newtonian when the flux-input of the moving matter to the gravitational field is bypassed for one reason or another (see § 8.3 and § 8.4 below). Therefore, incorporating the peculiar-flux input to gravity is what distinguishes the relativistic treatments from the rest and in so doing it can also provide a solution to the bulk-flow puzzle.

### v3

@@ 2.4.2 Resolution of historical conflicts
Another significant resolution concerns the apparent conflict between peculiar velocity surveys and other velocity-sensitive probes. An analysis by Boruah et al [159] showed that, once sparse sampling, survey selection, and cosmic variance are properly accounted for within consistent modeling frameworks, independent probes of large-scale motions yield mutually consistent constraints on the velocity field. Furthermore, it has been argued (see [21]) that properly accounting for the survey window function and sparse sampling can reconcile the apparent conflict between bulk flow measurements and CDM predictions, attributing many discrepancies to the specific realization of structure within our local volume.
@@ 3.1 Fundamental Principles of Peculiar Velocity Measurement
Two primary methodological approaches have emerged for characterizing bulk flows. The first involves measuring individual galaxy peculiar velocities and then calculating their weighted average within specified volumes [26, 27]. The second employs likelihood methods to fit parametric models directly to the observed velocity field [7, 11].
@@ 3.3.2 Novel techniques and emerging methods
where is the Hubble constant, and is the speed of light. We note that this relationship is valid in the low-redshift limit (i.e. for ).
@@ 4 Theoretical framework
This section provides the basic theoretical background of the 1+3 covariant approach to relativistic cosmology, which will be used in the subsequent sections to study the evolution of the observed large-scale peculiar motions. In the process, the relativistic results will also be directly compared to those of the available Newtonian and quasi-Newtonian studies.
@@ 4.1 Covariant spacetime splitting
The covariant formalism dates back to the 1950s (e.g. see [211, 212]) and has since been used in numerous studies by many authors (see [99, 100] for recent reviews with more details and references). The approach utilises a local 1+3 splitting of the spacetime into time and 3-dimensional space. This is achieved by introducing a group of observers, “living” along timelike worldlines with their associated 3-spaces orthogonal to them. Instead of using the metric tensor, the covariant formalism employs the Ricci and the Bianchi identities, together with the Einstein field equations [99, 100].
Consider a group of observers with 4-velocity , where are their (generalised) coordinates and their proper time, so that . The -field determines the observers’ temporal direction and introduces a local 1+3 “threading” of the spacetime into time and 3-D space. Then, the tensor projects into the observers’ instantaneous rest space and also acts as the 3-D metric. Note that the projector is symmetric, spacelike and also satisfies the constraints and .
@@
Note that we adopt spacetimes with metric () of signature () and use geometrised units with , unless stated otherwise. Also, Latin indices take values between and , while their Greek counterparts run from to .
@@ 4.1.1 The gravitational field
respectively. In the above, is the totally antisymmetric Levi-Civita tensor of the observer’s 3-dimensional rest-space, with by construction and with representing its 4-dimensional analogue. Also note that has a Newtonian analogue but has not (e.g. see [213]). For this reason the electric Weyl tensor is primarily associated with the tidal field and the magnetic with gravitational waves, although both Weyl tensors are necessary for the propagation of gravitational waves. In addition, both and are spacelike, symmetric and trace-free tensors.
@@ 4.2.1 Kinematics
Here, is the volume scalar, is the shear, is the vorticity and is the 4-acceleration. The last three are purely spacelike, since . The volume scalar describes expansion/contraction when positive/negative respectively. In the homogeneous and isotropic FLRW models, is used to define the cosmological scale factor () by means of . The same relation also defines the mean scale factor in the anisotropic Bianchi cosmologies. Given its trace-free nature, the shear monitors shape distortions under constant volume, while antisymmetric vorticity tensor describes rotation and also defines the rotational axis via the vorticity vector . Finally, nonzero 4-acceleration indicates the presence of non-gravitational forces and ensures that the observers’ worldlines are not geodesics.
@@ 4.3 The Friedmann universes
Our discussion so far refers to cosmological models that are inhomogeneous and anisotropic. Nevertheless, observations and our theoretical prejudice (the CMB and the Copernican principle) support the homogeneous and isotropic Friedmann-Robertson-Walker (FLRW) models.
4.3.1 The FLRW metric
@@
4.3.2 FLRW cosmologies
The high symmetry of the Friedmann models ensures that time-dependent scalars are the only surviving variables. Hence, in covariant terms, the FLRW cosmologies have , where is the familiar Hubble parameter, and . In addition, the isotropy of the Friedmann universes ensures that they can only accommodate perfect fluids, with and ). Finally, due to spatial homogeneity, all the 3-gradients (e.g. , , etc) vanish by default. All this means that the only nontrivial equations are the FLRW versions of Raychaudhuri’s formula, together with the continuity and the Friedmann equations. These follow from (4.2.3), (4.2.11a) and (4.2.16) and are given by
@@ 4.3.2 FLRW cosmologies
In FLRW universes with conventional matter, effectively coincides with the particle horizon (i.e. ), in which case the Hubble radius determines the regions of causal contact as well.
@@
with in the open model and in the closed. Consequently, always for , with as and when . On the other hand, for and when in closed () models.
@@ 4.3.3 Luminosity distance
The luminosity distance of a source at redshift is , with and being the present values of the scale factor and of the source’s radial distance (e.g. see [220]). The latter follows by integrating the line element of a null geodesic. In a spatially flat FLRW model, the integration gives
@@
It should be noted that the above analysis applies to exact FLRW models free of peculiar-velocity perturbations, or of any other type of distortions for that matter. Alternative expressions of the luminosity distance, as well as generalisations of that incorporate the (linear) effects of peculiar motions, can be found in § 10.1.1.
4.3.4 Scale-factor evolution in FLRW cosmologies
In the following sections, the host universe will typically be a spatially flat FLRW model with . Nevertheless, there will also be cases where the background spacetime will be allowed to have nonzero spatial curvature (see § 6.3.6 and § 7.1.8 below). It will therefore help to briefly go through the basic evolution laws of Friedmann universes with as well.
@@ 4.3.4 Scale-factor evolution in FLRW cosmologies
The case: In FLRW spacetimes with curved spatial sections it helps to parametrise the solutions in terms of the conformal time ( – defined so that ). Then, for , and relations (4.3.3) and (4.3.4) combine to the parametric solution
@@ 5 Cosmologies with a peculiar-velocity “tilt”
By default, the study of peculiar motions requires tilted cosmologies, equipped with two (at least) families of observers/frames moving relative to each other. Then, by construction, at every spacetime event there are two coordinate systems. This in turn guarantees that boosting from one frame to the other does not violate causality.
@@ 5.1 Relatively moving observers
Peculiar velocities are typically defined and measured with respect to the reference frame of the universe, which has been largely identified with that of the cosmic microwave photons. By definition, the latter is the only coordinate system where the CMB dipole vanishes. While only idealised observers follow the cosmic microwave frame, all those living in typical galaxies (like our Milky Way) are expected to have some finite peculiar velocity with respect to the CMB. Typically, these realistic observers measure their own peculiar velocity by the magnitude and the direction of the observed CMB dipole. This way, for example, the peculiar velocity of our Local Group of galaxies has been estimated at approximately 600 km/sec (e.g. see [224, 225]). The reader is also referred to [226] for further discussion and references.
Let us consider a tilted spacetime filled with a multi-component fluid and allow for several 4-velocity fields in relative motion at every spacetime event. Suppose also that is the reference 4-velocity of the CMB frame and that of the -th fluid component (see Fg. 1), with . Then, the symmetric 3-tensors
project orthogonal to and respectively. By construction, the above projectors satisfy the constraints , , and .
@@
One can recast the velocity boost (5.1.4) in terms of the hyperbolic “tilt” angle () between the two 4-velocity vectors (see Fig. 1). Indeed, given that and that (see Fig. 1 above and also [227]), we have
@@ 5.3 Linear relations between the frames
Consider an FLRW background universe, where the only non-vanishing variables are scalars that depend solely on time. These are the density (), the isotropic pressure (), the volume expansion (, with ) and the spatial Ricci scalar (, with and – see § 4.3.2 earlier). Then, when dealing with non-relativistic peculiar velocities (with and ), the nonlinear relations (5.2.4)-(5.2.7) linearise to
@@
respectively. Recall that is the peculiar velocity of the tilted frame relative to that of the CMB (see Fig. 1 in § 5.1 before). Then, is the local volume scalar, with since the peculiar flow can expand or contract locally, is the local votricity vector and is the local shear tensor of the peculiar motion. The above follow from the spatial gradient of the peculiar-velocity field, which decomposes as [228]
@@ 6 Linear peculiar velocities
Bulk peculiar flows are commonplace in the universe, confirmed by several surveys over many years (e.g. see [229] for the Planck 2013/2015 constraints). Such large-scale motions are believed to have started as weak peculiar-velocity perturbations around the time of recombination, which then grew to the observed sizes and speeds driven by structure formation. There are open issues, however. The reported velocities generally lie within CDM expectations on relatively small scales, roughly Mpc [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23], but appear in excess of the typical CDM predictions on larger scales [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34].555The Planck constraints disfavour the so called “dark flows”. The latter are extreme bulk flows, which sizes and speeds well in excess of the current CDM limits (see § 9.4 below). In this section we will review the theoretical studies of peculiar velocities, referring the reader to § 9 for the observational surveys and reports.
@@ 6.1.1 Newtonian peculiar kinematics
In what follows, we will look at the Newtonian covariant approach to the linear evolution of peculiar velocities in cosmology during the Einstein-de Sitter epoch of the universe. This will also facilitate the direct comparison of the purely Newtonian results with those of the subsequent quasi-Newtonian and relativistic studies (see § 6.2 and § 6.3 respectively). The linear nature of all these treatments confines them to sufficiently large scales, which typically lie close and beyond the 100 Mpc threshold. For a discussion on the available nonlinear studies, the reader is referred to § 8.4 - § 8.6.
@@
where is the peculiar velocity of the -field relative to the (reference) -frame.666Typical Newtonian studies of peculiar motions use physical () and comoving () coordinates, with . The time derivative of the latter leads to , where , and are the total, the Hubble and the peculiar velocities respectively. On an FLRW background, the above velocity relation is equivalent to the one seen in Eq. (6.1.1). The irreducible kinematics of the -field are given by the decomposition
@@
with , and respectively representing the volume scalar, the shear tensor and the vorticity of the peculiar flow.777Comparing (6.1.2) to its relativistic counterpart (see decomposition (4.2.1) in § 4.2.1), one notices the absence of an acceleration term from the right-hand sides of the Newtonian expression. This reflects the fact that, in Newtonian physics, space and time are separate and independent entities. As before (see Eq. (5.3.3) in § 5.3), positive/negative values for indicate (locally) expanding/contracting peculiar flows. Also, and are respectively the shear and vorticity of the peculiar-velocity field. Starting from (6.1.1), one arrives at the following (nonlinear) relations
@@ 6.1.2 Linear sources of peculiar velocities
In a perturbed almost-FLRW Newtonian universe (filled with a pressureless perfect fluid – baryonic or/and CDM – with ), the peculiar-velocity field is a linear perturbation. Then, by setting in the matter frame, expression (6.1.5) recasts to
Hence, the driving force of linear peculiar velocities is the acceleration, which for zero pressure is given by (with being the gravitational potential). As a result, Eq. (6.1.6) reads
@@ 6.2.1 The quasi-Newtonian setup
The quasi-Newtonian approach is used to obtain Newtonian-like equations and solutions, by adopting a reference frame with zero linear vorticity and shear (i.e. at the linear level by default). This, however, imposes strict restrictions on the spacetime and compromises its relativistic nature. For instance, in the quasi-Newtonian models, the magnetic component of the Weyl field also vanishes, which in turn means that gravitational waves are switched off as well. Perturbed relativistic spacetimes without these features are rather unnatural and the problematic nature of the quasi-Newtonian approach has been noted in [100] (see § 6.8.2 there for related discussion and “warning” comments). One should therefore be wary before adopting the quasi-Newtonian approximation in relativistic cosmological studies.
When it comes to the study of linear peculiar velocities, the downside of using the quasi-Newtonian approach is not the absence of gravitational waves, but the fact that the (purely relativistic) gravitational input of the peculiar flux (see § 5.3 earlier) is not accounted for. This is an additional side-effect of the zero vorticiy and shear constraints. The direct consequence is that the driving force of the linear peculiar-velocity field, namely the 4-acceleration, is given by the gradient of an ad hoc (Newtonian-like) scalar potential. Moreover, the potential has no expression to describe it and its evolution follows from an ansatz that is not uniquely determined. All this may simplify the calculations, but at the same time it seriously compromises the relativistic nature of the treatment and blurs the physics. As a result, the quasi-Newtonian analysis (inadvertently) reduces to Newtonian, which means that it cannot explain the fast and deep bulk peculiar flows reported in recent surveys, like those of [33, 34] for example. In contrast, as will show in § 6.3 below, the relativistic analysis provides theoretical support to the aforementioned bulk flows, without imposing any constraints upon the perturbed spacetime. Al this makes the related quasi-Newtonian studies an interesting mathematical exercise but without any physical input.
@@ 6.2.2 Linear quasi-Newtonian approach
The starting point is a tilted, perturbed Einstein-de Sitter background universe, with two families of observers in relative motion. These are the idealised (reference) CMB observers with 4-velocity and their realistic (tilted) counterparts with 4-velocity . The latter are assumed to move with (non-relativistic) peculiar velocity with respect to the former, so that (see Fig. 2).888In the background spacetime, the peculiar-velocity field vanishes by default and the two 4-velocities coincide. This makes a gauge-invariant linear perturbation [231]. Assuming that the pressureless matter follows timelike geodesics, one may set in the tilted frame. Then, in line with (5.3.1b) and (5.3.4c), we have
@@
There is a serious downside, however, because is an ad hoc potential with no expression for it. In addition, the time evolution of the quasi-Newtonian potential is not uniquely determined and follows from the ansatz , which requires setting the shear to zero as well [42]. Moreover, when it comes to the study of peculiar velocities, the key disadvantage is that the 4-acceleration adopted in Eq. (6.2.2) does not account for the gravitational input of the peculiar flux and this severely compromises the relativistic nature of the analysis. After all this, the linear evolution of the quasi-Newtonian 4-acceleration is given by [42, 43]
@@
Before closing this section, we should point out that the general problems of the quasi-Newtonian studies, namely of adopting a reference frame with zero linear shear and vorticity (and the rest of the constraints that follow), have been known and they are clearly stated in § 6.8.2 of [100]. However, the extent and the depth of the problem was not realised because there was no direct comparison with a proper relativistic treatment that did not employ the quasi-Newtonian frame. Next, we will provide such a comparison by avoiding the quasi-Newtonian zero vorticity and shear restrictions, when studying linear peculiar velocities. In the process, we will also identify the pivotal role of the peculiar flux and show how deeply damaging it can be to bypass it. Without accounting for the gravitational input of the flux, peculiar-motion studies that may appear relativistic reduce to Newtonian.999In the literature there additional examples of studies involving peculiar motions, which start relativistically but reduce to Newtonian when the flux-input of the moving matter to the gravitational field is bypassed for one reason or another (see § 8.3 and § 8.4 below). Therefore, incorporating the peculiar-flux input to gravity is what distinguishes the relativistic treatments from the rest and in so doing it can also provide a solution to the bulk-flow puzzle.